# Power allocation for cell-free MIMO integrated sensing and communication

### v1

@@ IV Problem Formulation
where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding performance thresholds  [^3] The thresholds can be set by referring to the CRLBs under a uniform power allocation..
Similar to the approaches in [40, 49, 11, 13], we assume that the target’s RCS, location, and velocity have been coarsely estimated by the ISAC system and are periodically updated over time [^4] The previously estimated target parameters can be reused for PA optimization at the current time step, particularly in consecutive estimation or target tracking scenarios.. Therefore, the PA coefficient vector $\bm{\rho}$ becomes the sole optimization variable, while the other relevant parameters are implicitly embedded through their coupling in equations (22) and (25).
V Algorithm Derivation
In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (57a)-(26e). However, the strong nonlinearity and non-convexity of the CRLB-based constraints in (26d) and (26e) render the problem analytically intractable. To overcome this challenge, we adopt a conjugate gradient (CG) iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
We first introduce a penalty (function) for violating the nonlinear inequality constraints in (26c)-(26e)
$\alpha({\bm{\rho}})$ $=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (27)
@@ V Algorithm Derivation
$\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (28)
with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$. We consider $q=2$ to guarantee the continuous differentiable penalty functions. Then, the objective function for the penalty problem can be written as
$L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}).$ (29)
As mentioned in [54], as the penalty factor $\mu$ is made large, the points generated for minimizing (29) approach an optimal solution of the original objective function (57a) from outside the feasible region and the iteration process terminate with $\mu\alpha({\bm{\rho}})\leq\epsilon_{\mu}$ for a small threshold $\epsilon_{\mu}>0$. This issue can be solved by setting $\epsilon_{\mu}$ small enough or employing much tighter thresholds than the required bounds to some extent to mitigate the loss due to the non-zero threshold $\epsilon_{\mu}$.
On the other hand, intuitively, when orthogonal waveforms are employed, increasing the transmission power—or equivalently, the signal-to-noise ratio (SNR)—at any individual transmitter enhances both sensing and communication performance. Therefore, it is desirable to fully utilize the total power budget, as specified in (26b). To enforce this equality constraint and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function
${\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (30)
where ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$. It is evident that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ holds with ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. Thus, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of any $N\times 1$ vector ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$. According to (30), the objective function can be further written as,
$\mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (31)
This objective function makes all the constraints satisfied and thus convert the original constrained optimization regarding ${\bm{\rho}}$ into an unconstrained one regarding ${\bm{\varpi}}$.
V-A Modified Conjugate Gradient Algorithm
We now propose a projection based modified CG (MCG) algorithm to minimize (31). Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
(ϖ)=bold-italic-$\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]}.$ (32)
where
$\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
$+\sum_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big{)},$ (33)
with components
$\begin{cases}\nabla\tilde{f}_{l}({\bm{\rho}})=2\max\{f_{l}({\bm{\rho}}),0\}\nabla f_{l}({\bm{\rho}})\\ \nabla\tilde{f}_{v}({\bm{\rho}})=2\max\{f_{v}({\bm{\rho}}),0\}\nabla f_{v}({\bm{\rho}})\\ \nabla f_{\min,n}({\bm{\rho}})=2\min\{\rho_{n}-\rho_{\min,n},0\}{\bm{e}}_{n}\\ \nabla f_{\max,n}({\bm{\rho}})=2\max\{\rho_{n}-\rho_{\max,n},0\}{\bm{e}}_{n}.\end{cases}$ (34)
@@ V-A Modified Conjugate Gradient Algorithm
Note that ${\bm{e}}_{n}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
Solving (31) using an iterative method yields the update
${\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (35)
where ${\bm{d}}_{i}$ is the descent direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The descent direction using the MCG is updated by
${\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (36)
where $\varsigma_{i}$ denotes the deflection factor that characterizes a particular CG algorithm. Note that $-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$ is still the descent direction of $L({\bm{\rho}}_{i})$ with $-\nabla L({\bm{\rho}}_{i})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix.
@@
Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$. Then, taking the projection of (35) leads to
${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}.$ (37)
which has the same form as (35). Therefore, we hereafter directly adopt the update in (37). Moreover, note that the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$ holds for all $i>1$ if and only if ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$. This property follows from the iteration rules in (36) and (37).
To enable the convergence, it is necessary to develop a suitable step size $\upsilon_{i}$ and search direction ${\bm{d}}_{i}$ in each iteration to guarantee the non-increasing objective function i.e.,
$L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (38)
Firstly, for the non-increasing search direction for each iteration $i$, it requires
${\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (39)
where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. Using (36), we have,
$\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (40)
where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$
$\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (41)
@@
Note that $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
@@
When considering inexact line searches, Hestenes and Stiefel’s [55, 56] choice for $\varsigma_{i}$ can be preferable. We have the modified HS deflection,
$\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (42)
@@
which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$. In fact, when using the MCG algorithm, the orthogonality is desired between the (conjugate) gradients in adjacent iterations, i.e.,
@@
Equations (43) and (44) ensure the positivity of both ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively. This allow for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$. Furthermore, (43) and (44) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively. As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (40).
@@
Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we will also let $\varsigma_{i}=0$ to restart the new loop of MCG iteration. A proper reboot is potential in mitigating the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$. Then, we have $f_{i}(0)=L({\bm{\rho}}_{i})$. The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
$f_{i}(\upsilon_{i})$ $\thickapprox f_{i}(0)+\upsilon_{i}\nabla f_{i}(0)+\upsilon_{i}^{2}\nabla^{2}f_{i}(0)/2$

### v2

@@ III-B Problem Formulation
where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds.
IV Power Allocation Algorithm for CF MIMO-ISAC
In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (22a)-(22e). Unfortunately, the strong nonlinearity and non-convexity of the CRLB-based constraints in (22d) and (22e) render the problem analytically intractable. To overcome this challenge, we adopt a CG iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
To derive a numerical solution, we start from converting the original constrained optimization problem into an unconstrained one. To this end, we first introduce a penalty (function) for violating the nonlinear inequality constraints in (22c)-(22e), defined as [39]
$\alpha({\bm{\rho}})=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum\nolimits_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$
where the involved penalty sub-functions are given respectively by
$\small\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (23)
with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$. $q$ is the power of the penalty function. We consider $q=2$ to guarantee their continuous differentiability. Then, the objective function for the penalty problem can be written as
$\small L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}),$ (24)
where $\mu$ denotes the penalty factor.
Additionally, to enforce this equality constraint (22b) and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function, i.e.,
$\small{\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (25)
where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$. We have ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds by initializing ${\bm{\rho}}_{0}$ such that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$.
By substituting (25) into (24), we can convert the original constrained optimization to an unconstrained optimization problem as follows:
$\small\arg\min\limits_{{\bm{\rho}}}$ $\ \mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (26)
Then, unconstrained optimization methods, e.g., CG and SD methods, can be employed to search for the optimal solution.
IV-A Modified Conjugate Gradient Algorithm
To numerically solve (26), we propose a projection-based MCG algorithm. The update at the $i$th iteration is given by
$\small{\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (27)
where ${\bm{d}}_{i}$ is the search direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The search direction using the MCG is updated by
${\bm{d}}_{i+1}$ $=-\nabla\mathcal{L}({\bm{\varpi}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (28)
where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
(ϖ)=bold-italic-$\small\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]},$ (29)
@@ IV-A Modified Conjugate Gradient Algorithm
$\small\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
$+\sum\nolimits_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big{)},$ (30)
and
$\small\begin{cases}\nabla\tilde{f}_{l}({\bm{\rho}})=2\max\{f_{l}({\bm{\rho}}),0\}\nabla f_{l}({\bm{\rho}})\\ \nabla\tilde{f}_{v}({\bm{\rho}})=2\max\{f_{v}({\bm{\rho}}),0\}\nabla f_{v}({\bm{\rho}})\\ \nabla f_{\min,n}({\bm{\rho}})=2\min\{\rho_{n}-\rho_{\min,n},0\}{\bm{e}}_{n}\\ \nabla f_{\max,n}({\bm{\rho}})=2\max\{\rho_{n}-\rho_{\max,n},0\}{\bm{e}}_{n}.\end{cases}$ (31)
@@
Note that ${\bm{e}}_{n}\in\mathbb{R}^{N\times 1}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
@@
Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$. Thus, taking the projection of (27) leads to
$\small{\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (32)
where the search direction is updated by
$\small{\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla\mathcal{L}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}.$ (33)
We have ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>1}$ due to ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{1}=0$ and ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Therefore, the equality constraint in (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at $i$th iteration for $\forall{i>1}$ if and only if the initial PA vector ${\bm{\rho}}_{1}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$.
We now investigate the selection of the deflection factor $\varsigma_{i}$ and step size $\upsilon_{i}$ at each iteration, which is significant in guaranteeing the convergence of the MCG algorithm. One necessary condition for the convergence is the non-increasing objective function at each iteration, i.e.,
$\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (34)
To this end, the search direction ${\bm{d}}_{i}$ should be the descent direction and the step size $\upsilon_{i}$ should be correspondingly upper bounded by a threshold at each iteration.
IV-A1 Descent direction
To ensure a descent direction ${\bm{d}}_{i+1}$ at the $(i+1)$th iteration, it requires
$\small{\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (35)
where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. The initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$ is the descent direction of $L({\bm{\rho}}_{1})$ with $-\nabla L({\bm{\rho}}_{1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix. Substituting (33) into (35), we have
$\small\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (36)
where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$ with its expression given by
$\small\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (37)
@@ IV-A1 Descent direction
The numerator satisfies $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
@@
When employing inexact line searches, the deflection selection proposed by Hestenes and Stiefel [40], referred to as HS method, for $\varsigma_{i}$ is preferable. Accordingly, we adopt the modified HS deflection, given by
$\small\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
@@
which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$. In fact, when using the MCG algorithm, it is desirable to maintain orthogonality between the conjugate gradients of adjacent iterations, i.e.,
@@
(39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively. This allows for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$. Furthermore, (39) and (40) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively. As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (36).
@@
Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations. Such a reboot strategy can help mitigate the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
IV-A2 Step size optimization
To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$. The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
$\small f_{i}(\upsilon_{i})$ $\thickapprox f_{i}(0)+\upsilon_{i}\nabla f_{i}(0)+\upsilon_{i}^{2}\nabla^{2}f_{i}(0)/2$